OpenStudy (awesomedude2078):

Calculate the total paid hours, and the weekly pay at $5.15 per hour, of the following workers (multiply pay by 1.5 for overtime hours). Regular Hours Overtime Hours Weekly Pay Tim Carter, who worked 50 hours. 40 ? ? Jesse Jones, who worked 47½ hours. 40 ? ? Barbara Burns, who worked 44 hours. 40 ? ?

563blackghost (563blackghost):

Still need help with this?

563blackghost (563blackghost):

I'll show you on how to do the first one.... We know Tim Carter worked a total of `50 hours for the week` but he had only `40 regular hours` so we would multiply `40 by the weekly pay of $5.15 per hour` So... \(\huge{40 \times 5.15=$206}\) Now he worked `10 extra hours` that would mean we would multiply that by weekly pay and then multiply it by `1.5`.... \(\huge{(5.15 \times 10)\times 1.5 = $77.25}\) Next you would add the sum of the two equations to find how much money he got for the week... \(\huge{$206+$77.25=$283.25}\)